OpenStudy (anonymous):

calc 3 help use the double integral to find the volume of the solid in the first octant bounded above by paraboloid z=x^2+3y^2, below by the plane z=0, laterally by y=x^2 and y=x

OpenStudy (anonymous):

I simply used o to 1 for dx and x^2 to x for dy, integrated and got 11/70 can someone check my answer? That would be awesome

OpenStudy (anonymous):

\[ \int _0^1\int _{x^2}^x\left(x^2+3 y^2\right)dydx=\frac{11}{70 } \]

OpenStudy (anonymous):

Your answer is right.